LUYVX:

Im the table, note your guesses and checks to find the length of tile B, which has an area of 10 square inches

foxy44:

Umm do you have the answer, like what might be possible?

foxy44:

Like multiple choice?

LUYVX:

No,

1 attachment
foxy44:

hmm.....I don't think i can help but i think i can get someone to help you.

foxy44:

umm sorry for the wait just one moment.

foxy44:

umm did you answer yet or get any help?

foxy44:

foxy44 6 minutes ago umm someone needs help ASTRO!! i'm losing a patient here. lol Astro 4 minutes ago sorry foxy44 4 minutes ago LUYVX. closed posts Astro 2 minutes ago i dont feel like it foxy44 a minute ago :O Ah man, I tried so hard to help.

LUYVX:

Thanks for trying tho

SmokeyBrown:

Hi, and welcome to QuestionCove! To start, we know the (square?) tile has an area of 10 square inches. If the tile is a square, then the length of one side should be the "square root" of the area. In other words, if we take the length of one side and multiply it by itself, we will get the area. That is a hint about how we should make our guesses. We know that 9 is smaller than 10, and the square root of 9 is 3 because 3*3=9. So, the length of tile B must be larger than 3 but smaller than 4, since 4*4=16, which is larger than 10. We can also probably say that the length of tile B is closer to 3 than it is to 4, since 9 is closer to 10 than 16 is to 10. Maybe we make our first guess to be 3.1; in this case, we can check: 3.1*3.1 = 9.61 Which is less than 10. So, our next guess should be a bit bigger. We'll keep on making guesses and checking them until we get as close to 10 as we can. I made the first guess for you; would you like to continue making the remaining 5 guesses and checks?

SmokeyBrown:

*remaining 4 guesses, excuse me :)